# Philadelphia, PA, USA

### v1

@@ I Introduction
In this paper, we define a constrained learning framework to enable dynamic network slicing in Wi-Fi. Sections II and III introduce a flexible Wi-Fi network slicing framework and the problem setup. Departing from the aforementioned works, we formulate the network slicing problem as a constrained radio resource management (RRM) optimization problem and develop an unsupervised state-augmented primal-dual algorithm in Section IV. The notion of state augmentation—proposed in [11] for constrained RL and in [12] for wireless resource allocation—incorporates the Lagrangian dual multipliers into the state space similar to a closed-loop feedback system. This results in a practical online algorithm that samples from an optimal policy and exhibits feasibility and near-optimality guarantees, which do not hold for the regularized and conventional primal-dual methods. We demonstrate these attributes of state augmentation in the context of network slicing with ergodic throughput and latency QoS requirements through numerical experiments in Section V.
@@ III Network Slicing Model
The network operates over $T$ discrete epochs, indexed by $t=0,\ldots,T-1$. We refer to the discrete epoch $t$ as the $t$th slicing window. Each slicing window lasts $\tau_{\max}$ seconds and we index any timestamp within that slicing window with $\tau\in[0,\tau_{\max})$. Each flow $i$ maintains a queue of size $Q^{t}_{i}(\tau)\in[0,Q_{\max}]$ that evolves with a dynamic traffic arrival rate $\alpha^{t}_{i}(\tau)\sim\alpha(\bar{\mu}^{t}_{i})$ and a dynamic transmission rate $r^{t}_{i}(\tau)$.
@@
The QoS metrics throughput and latency for a flow $i\in{\mathcal{U}}$ over the $t$th slicing window, denoted by $\bar{r}_{i}(t)$ and ¯$\bar{\ell}_{i}(t)$, respectively, are determined primarily by the dynamic rate $r^{t}_{i}(\tau)$ which itself is determined by the frequency resource units (FRUs) allocated to flow $i$ by the AP and the channel state which is an indicator of the signal-to-interference-plus-noise ratio (SINR) experienced by the device receiving flow $i$. The channel state varies randomly from each slicing window to window, e.g., following a Rayleigh fading channel model.
@@
that takes the network state ${\mathbf{s}}_{t}$ as input and outputs the corresponding slicing decisions. Our goal is to optimize the parameter vector $\boldsymbol{\theta}$ to maximize the system performance subject to the QoS requirements and the total bandwidth budget of $p_{\mathrm{H}}(t)+p_{\mathrm{L}}(t)+p_{\mathrm{B}}(t)\leq 1$. Note that we can drop the total bandwidth budget constraint in subsequent derivations since it holds with equality for any optimal slicing decision and can be implicitly satisfied by the design of the learning architecture (e.g., a softmax layer).
@@
where $f_{0}:\mathcal{S}\times\mathbb{R}^{3}\to\mathbb{R}$ is the objective (e.g., negated total network throughput), ${\mathbf{f}}:\mathcal{S}\times\mathbb{R}^{3}\to\mathbb{R}^{c}$ is the vector-valued constraint function (e.g., minimum throughput and maximum packet latency QoS) for $c$ constraints in total.
@@ IV Proposed State-Augmented Algorithm
The training and test phases of the state-augmented algorithm are summarized in Fig. 2. We note that under some mild technical assumptions, near-optimality and almost sure feasibility guarantees of (6) extend to the state-augmented primal-dual algorithm. Moreover, although we do not consider a CRL problem, our proposed algorithm is amenable to tackling CRL problems. We refer the reader to [11, 12] for a theoretical exposition on the state-augmented primal-dual algorithms.
@@ V Experimental Results
Given a FRU allocation ${\mathbf{p}}(t)$ and associated channel state ${\mathbf{h}}^{t}(\tau)$, we can define the average throughput $\bar{r}_{i}(t)$ over a slicing period $t$ for each flow $i$ as follows. The instantaneous data rate with a current channel state $h$ and corresponding SINR is given by the Shannon rate $g(h):=\log(1+h/\sigma^{2})$. The full instantaneous rate for flow $i$ can be defined as
@@
$\bar{r}_{i}(t)=\overline{r^{t}_{i}(\tau)}=\mathbb{E}_{\tau\in[0,\tau_{\max}]}\left[r^{t}_{i}(\tau)\cdot\beta^{t}_{i}(\tau)\right]$ (13)
@@
where $r^{t}_{i}(\tau_{j})$ is the instantaneous rate during transmission of packet $j$ as defined in (12), $P>0$ is the fixed packet size, $\alpha_{j}$ and $\tau_{j}$ are the arrival and transmission times of packet $j$, respectively. Then the packet delay is defined as the sum of queueing delay $\tau_{j}-\alpha_{j}$ plus the actual transmission delay incurred. We finally define the latency for a given flow over a slicing window as the maximum packet latency, i.e.,
¯$\bar{\ell}_{i}(t):=\max_{j\in\mathcal{J}^{t}_{i}}\ell^{t}_{i}(j).$ (15)
V-A Experimental Setup
Figure 3: Test evolutions of high-throughput (left), low-latency (middle-left) QoS constraints, best-effort throughput objective (middle-right) and average rate of QoS violations (right) for the proposed state-augmented primal-dual algorithm and the baselines. Objective and the constraints are averaged (solid lines) over $128$ test network realizations and the shadowed bands cover to the worst ($99^{\text{th}}$ percentile) realization (dashed lines). Red dashed lines indicate the respective QoS requirements.
The total bandwidth is $W=20$ MHz. We fix the total number of flows to $|{\mathcal{U}}|=20$ and randomly sample the (initial) number of H, L and B flows for each network configuration. Packet generations and traffic arrival for all flows follow a constant bit rate model, mean packet generation rates $\overline{\mu}^{t}_{i}$ are (uniform) randomly sampled and dynamic packet generation rates $\alpha(\overline{\mu}^{t}_{i})$ vary across slicing windows following a random-walk with noise standard deviation of $0.5$ bps/Hz. The initial average data rates are sampled randomly $\mathrm{Unif}[1,5]$ bps/Hz for high-throughput and best-effort flows and $\mathrm{Unif}[0.5,1.5]$ bps/Hz for the low-latency slices. We adopt the Rayleigh fading channel model and the transmitter and access-point (AP) deployment approach of [12].
@@ V-A Experimental Setup
$\max_{\boldsymbol{\phi}\in\boldsymbol{\Phi}}$ $\frac{1}{T}\sum_{t=0}^{T-1}\frac{1}{|\mathcal{U}_{\mathrm{B}}|}\sum_{i\in\mathcal{U}_{\mathrm{B}}}\bar{r}_{i}\big{(}{\mathbf{s}}_{t},{\mathbf{p}}^{\boldsymbol{\phi}}({\mathbf{s}}_{t},\boldsymbol{\lambda}_{\lfloor t/T_{0}\rfloor};\boldsymbol{\phi})\big{)}$ (16)
$\mathrm{s.t.}$ $\frac{1}{T}\sum_{t=0}^{T-1}\left[\max_{i\in\mathcal{U}_{\mathrm{H}}}\left(1-\frac{\bar{r}_{i}\big{(}{\mathbf{s}}_{t},{\mathbf{p}}^{\boldsymbol{\phi}}({\mathbf{s}}_{t},\boldsymbol{\lambda}_{\lfloor t/T_{0}\rfloor};\boldsymbol{\phi})\big{)}}{r_{\min}}\right)\right]\leq 0,$
$\frac{1}{T}\sum_{t=0}^{T-1}\left[\max_{i\in\mathcal{U}_{\mathrm{L}}}\left(\frac{\bar{\ell}_{i}\big{(}{\mathbf{s}}_{t},{\mathbf{p}}^{\boldsymbol{\phi}}({\mathbf{s}}_{t},\boldsymbol{\lambda}_{\lfloor t/T_{0}\rfloor};\boldsymbol{\phi})\big{)}}{\ell_{\max}}-1\right)\right]\leq 0.$
The objective is to maximize the average throughput for best-effort slices. We formulate the QoS constraints as the ergodic average of the worst-case metrics over the respective SLAs to account for varying number of flows.
@@
V-B Simulation Results & Discussion
In Fig. 3, we present the test performance of the primal-dual (PD) and state-augmented primal-dual (SA-PD) algorithms against the baselines. The first three subfigures from the left show the evolution of the test performance of the algorithms with respect to the training epochs. We validate that the state-augmented algorithm meets the throughput and latency QoS requirements while none of the traditional baselines meet the constraints simultaneously. The primal-dual algorithm iterates, although feasible for an average network, do not converge to strictly feasible policies for all networks and constraints simultaneously unlike the state-augmented algorithm. In the righmost subfigure, we report the average rate of instantaneous QoS violations and the state-augmented algorithm achieves a low rate of instantaneous violation for both constraints on top of ergodic feasibility. Furthermore, we plot the evolution of the QoS constraints for the worst network realization across slicing windows for fully trained models in Fig. 4 and verify the ergodic feasibility of the state-augmented slicing policy.
Figure 4: Time evolution of cumulative ergodic averages of the QoS constraints for the worst realization.
@@ V-B Simulation Results & Discussion
Figure 6: Example of policy switching with state-augmentation.
@@
In Fig. 5, we illustrate the online execution of the state-augmented algorithm for an example test network. As the throughput constraint is initially violated, the corresponding dual multipliers and the slice weights allocated to high-throughput slices increase until the throughput constraint is met. Fig. 6 exemplifies the policy switching behavior of the state-augmented algorithm where alternating spikes in dual multipliers lead to switching between prioritizing high-throughput and low-latency slices to maintain the feasibility of both constraints. Finally, we report in Table I the average rate of instantaneous and ergodic QoS violations for varying constraint level specifications. An ergodic throughput QoS violation occurs for a flow $i\in\mathcal{U}_{\mathrm{H}}$ whenever $(1/T)\sum_{t=0}^{T-1}\bar{r}_{i}(t)<r_{\min}$ happens. Ergodic latency violations are defined similarly. Both primal-dual algorithms outperform the traditional baselines and the state-augmented algorithm generally attains comparable, if not better, ergodic QoS satisfaction.
@@ VI Conclusion and Future Work
We formulated (Wi-Fi) network slicing as a constrained learning problem. Through numerical simulations, we illustrated that a state-augmented algorithm that minimizes an expected Lagrangian offline and updates the dual multipliers online lends itself to a practical primal-dual algorithm with feasibility and optimality guarantees for network slicing. Extension to a state-augmented RL algorithm for network slicing and end-to-end evaluation on a more realistic simulation platform are interesting future research directions.

### v2

@@ I Introduction
In this paper, we define a constrained learning framework to enable dynamic network slicing in Wi-Fi. Sections II and III introduce a flexible Wi-Fi network slicing framework and the problem setup. Departing from the aforementioned works, we formulate the network slicing problem as a constrained radio resource management (RRM) optimization problem and develop an unsupervised state-augmented primal-dual algorithm in Section IV. The notion of state augmentation—proposed in [11] for constrained RL and in [12] for wireless resource allocation—incorporates the Lagrangian dual multipliers into the state space similar to a closed-loop feedback system. The novelty of our solution lies in that state-augmentation allows for training a single model to learn Lagrangian minimizing policies over a range of dual multipliers and running dual updates online to sample from different learned policies. This approach results in a practical online algorithm that samples from an optimal (stochastic) policy and exhibits feasibility and near-optimality guarantees, which do not hold for the regularized and conventional primal-dual methods. We demonstrate these attributes of state augmentation in the context of network slicing with ergodic throughput and latency QoS requirements through numerical experiments in Section V.
@@ III Network Slicing Model
The network operates over $T$ discrete epochs, indexed by $t=0,\ldots,T-1$. We refer to the discrete epoch $t$ as the $t$th slicing window. Each slicing window lasts $\tau_{\max}$ seconds and we index any timestamp within that slicing window with $\tau\in[0,\tau_{\max})$. Each flow $i$ maintains a queue of size $Q^{t}_{i}(\tau)\in[0,Q_{\max}]$ that evolves with a dynamic traffic arrival rate $\alpha^{t}_{i}(\tau)\sim\alpha(\mu^{t}_{i})$ and a dynamic transmission rate $r^{t}_{i}(\tau)$.
@@
The QoS metrics throughput and latency for a flow $i\in{\mathcal{U}}$ over the $t$th slicing window, denoted by $r_{i}(t)$ and $\ell_{i}(t)$, respectively, are determined primarily by the dynamic rate $r^{t}_{i}(\tau)$ which itself is determined by the frequency resource units (FRUs) allocated to flow $i$ by the AP and the channel state which is an indicator of the signal-to-interference-plus-noise ratio (SINR) experienced by the device receiving flow $i$. The channel state varies randomly from each slicing window to window, e.g., following a Rayleigh fading channel model.
@@
that takes the network state ${\mathbf{s}}_{t}$ as input and outputs the corresponding slicing decisions. We include the parametrization $\boldsymbol{\theta}$ in the superscript to avoid confusion when we introduce a state-augmented parametrization in the next section. Our goal is to optimize the parameter vector $\boldsymbol{\theta}$ to maximize the system performance subject to the QoS requirements and the total bandwidth budget of $p_{\mathrm{H}}(t)+p_{\mathrm{L}}(t)+p_{\mathrm{B}}(t)\leq 1$. Note that we can drop the total bandwidth budget constraint in subsequent derivations since it holds with equality for any optimal slicing decision and can be implicitly satisfied by the design of the learning architecture (e.g., a softmax layer).
@@
where $f_{0}:\mathcal{S}\times\mathbb{R}^{3}\to\mathbb{R}$ is the objective (e.g., negated total network throughput), ${\mathbf{f}}:\mathcal{S}\times\mathbb{R}^{3}\to\mathbb{R}^{c}$ is the vector-valued constraint function (e.g., minimum throughput and maximum packet latency QoS) for $c$ constraints in total. Note that the objective and the constraints are derived based on the ergodic averages of the respective metrics.
@@ IV Proposed State-Augmented Algorithm
A suitable choice of the update window hyperparameter $T_{0}$ depends on the variance of the constraint function and implies a trade-off between obtaining an unbiased, low variance estimate of the gradients and the number of iterations needed for the dual multipliers to converge to their near-optimal values. Similarly, given the varying nature of the physical metric each constraint pertains to (e.g., rate measured in megabits whereas latency measured in milliseconds), different dual step sizes can be used to update the different components of the dual multiplier vector. For clarity of presentation, we adopted the scalar notation to denote the dual step size throughout the paper and applied a suitable normalization of the constraint function in our experiments (see Section V-B).
The training and test phases of the state-augmented algorithm are summarized in Fig. 2. We note that under some mild technical assumptions, near-optimality and almost sure feasibility guarantees of (6) extend to the state-augmented primal-dual algorithm. Moreover, although we do not consider a CRL problem, our proposed algorithm is amenable to tackling CRL problems. We refer the reader to [11] for a theoretical exposition on the state-augmented primal-dual algorithms.
@@ V Experimental Results
V-A QoS Evaluation Metrics
Recall that we operate over $T$ slicing windows indexed by $t\in[0,T-1]$ and we further index any timestamp within each slicing window with $\tau\in[0,\tau_{\max})$. Given a FRU allocation ${\mathbf{p}}(t)$ and associated channel state ${\mathbf{h}}^{t}(\tau)$, we can define the (ergodic) average throughput $r_{i}(t)$ over a slicing period $t$ for each flow $i$ as follows. The instantaneous data rate with a current channel state $h$ and corresponding SNR is given by the Shannon rate $g(h):=\log(1+h/\sigma^{2})$. The full instantaneous rate for flow $i$ can be defined as
@@ V-A QoS Evaluation Metrics
$r_{i}(t)=\mathbb{E}_{\tau\in[0,\tau_{\max}]}\left[r^{t}_{i}(\tau)\cdot\beta^{t}_{i}(\tau)\right]$ (13)
@@
where $r^{t}_{i}(\tau_{j})$ is the instantaneous rate during transmission of packet $j$ as defined in (12), $P>0$ is the fixed packet size, $\alpha_{j}$ and $\tau_{j}$ are the arrival and transmission times of packet $j$, respectively. Then the packet delay is defined as the sum of queueing delay $\tau_{j}-\alpha_{j}$ plus the actual transmission delay incurred. We finally define the latency for a given flow over a slicing window $t$ as the maximum packet latency, i.e.,
$\ell_{i}(t):=\max_{j\in\mathcal{J}^{t}_{i}}\ell^{t}_{i}(j).$ (15)
V-B Experimental Setup
Figure 3: Test evolutions of high-throughput (left), low-latency (middle) QoS constraints, best-effort throughput objective (right) for the proposed state-augmented primal-dual algorithm and the baselines. Objective and the constraints are averaged (solid lines) over $128$ test network realizations and the shadowed bands cover to the worst ($99^{\text{th}}$ percentile) realization (dashed lines). Red dashed lines indicate the respective QoS requirements.
The total bandwidth is $W=20$ MHz. We fix the total number of flows to $|{\mathcal{U}}|=20$ and randomly sample the (initial) number of H, L and B flows for each network configuration. Packet generations and traffic arrival for all flows follow a constant bit rate model, mean packet generation rates $\mu^{t}_{i}$ are (uniform) randomly sampled and dynamic packet generation rates $\alpha(\mu^{t}_{i})$ vary across slicing windows following a random-walk with noise standard deviation of $0.5$ bps/Hz. The initial average data rates are sampled randomly $\mathrm{Unif}[1,5]$ bps/Hz for high-throughput and best-effort flows and $\mathrm{Unif}[0.5,1.5]$ bps/Hz for the low-latency slices. We adopt the Rayleigh fading channel model and the transmitter and access-point (AP) deployment approach of [12].
@@ V-B Experimental Setup
$\max_{\boldsymbol{\phi}\in\boldsymbol{\Phi}}$ $\frac{1}{T}\sum_{t=0}^{T-1}\frac{1}{|\mathcal{U}_{\mathrm{B}}|}\sum_{i\in\mathcal{U}_{\mathrm{B}}}r_{i}\big{(}{\mathbf{s}}_{t},{\mathbf{p}}^{\boldsymbol{\phi}}({\mathbf{s}}_{t},\boldsymbol{\lambda}_{\lfloor t/T_{0}\rfloor};\boldsymbol{\phi})\big{)}$ (16)
$\mathrm{s.t.}$ $\frac{1}{T}\sum_{t=0}^{T-1}\left[\max_{i\in\mathcal{U}_{\mathrm{H}}}\left(1-\frac{r_{i}\big{(}{\mathbf{s}}_{t},{\mathbf{p}}^{\boldsymbol{\phi}}({\mathbf{s}}_{t},\boldsymbol{\lambda}_{\lfloor t/T_{0}\rfloor};\boldsymbol{\phi})\big{)}}{r_{\min}}\right)\right]\leq 0,$
$\frac{1}{T}\sum_{t=0}^{T-1}\left[\max_{i\in\mathcal{U}_{\mathrm{L}}}\left(\frac{{\ell}_{i}\big{(}{\mathbf{s}}_{t},{\mathbf{p}}^{\boldsymbol{\phi}}({\mathbf{s}}_{t},\boldsymbol{\lambda}_{\lfloor t/T_{0}\rfloor};\boldsymbol{\phi})\big{)}}{\ell_{\max}}-1\right)\right]\leq 0.$
The objective is to maximize the average throughput for best-effort slices. We formulate the QoS constraints as the ergodic average of the worst-case metrics over the respective SLAs to account for varying number of flows. We normalize the throughput and latency metrics so that the dynamic ranges of the two constraints are similar and choose a scalar step size for the dual updates. Following (3), we make the dependence of the QoS metrics on the network state and the policy explicit.
@@
(a) QoS constraints
(b) Rate of QoS violations
Figure 4: (a) Time evolution of cumulative ergodic averages of the QoS constraints for the worst realization and (b) bar plot of average rate of QoS violations.
@@
V-C Simulation Results
V-C1 Test Performance and Comparison with Baselines
In Fig. 3, we present the test performance of the primal-dual (PD) and state-augmented primal-dual (SA-PD) algorithms against the baselines. The evolution of the test performances of the primal-dual algorithms with respect to the training epochs are plotted and we validate that the state-augmented algorithm meets the throughput and latency QoS requirements while none of the traditional baselines meet the constraints simultaneously. The primal-dual algorithm iterates, although feasible for an average network, do not converge to strictly feasible policies for all networks and constraints simultaneously unlike the state-augmented algorithm. Indeed, the worst throughput and latency metrics for the primal-dual algorithm (dashed purple line) oscillate around the QoS requirements. That the average objective value is lower for the primal-dual algorithms is a consequence of relatively strict QoS requirements. Ultimately, the best-effort slices have the lowest priority as the name implies.
Figure 6: Example of policy switching with state-augmentation.
V-C2 Online Execution of the State-Augmented Algorithm
The ergodic feasibility of the state-augmented slicing policy is verified in Fig. 4(a) where we plot the evolution of the QoS constraints for the worst network realization across slicing windows for fully-trained models. In Fig. 4(b), we report the average rate of instantaneous QoS violations and the state-augmented algorithm achieves a low rate of instantaneous violation for both constraints on top of ergodic feasibility.
Fig. 5 illustrates the online execution of the state-augmented algorithm for an example test network. As the throughput constraint is initially violated, the corresponding dual multipliers and slice weights allocated to high-throughput slices increase until the throughput constraint is met.
V-C3 Policy Switching
Fig. 6 exemplifies the policy switching behavior of the state-augmented algorithm where alternating spikes in dual multipliers lead to switching between prioritizing high-throughput and low-latency slices to maintain the feasibility of both constraints. Neither of the policies that prioritize only either one of the two slices solve the constrained problem at hand. Finally, we report in Table I the average rate of instantaneous and ergodic QoS violations for varying constraint level specifications. An ergodic throughput QoS violation occurs for a flow $i\in\mathcal{U}_{\mathrm{H}}$ whenever $(1/T)\sum_{t=0}^{T-1}r_{i}(t)<r_{\min}$ happens. Ergodic latency violations are defined similarly. Both primal-dual algorithms outperform the traditional baselines and the state-augmented algorithm attains comparable, if not better, ergodic QoS satisfaction.
V-D Practical Implications & Discussion
We hold the view that state-augmentation provides a principled alternative to direct policy randomization/averaging by learning a single parametrization to minimize the Lagrangian for all dual multipliers and running dual dynamics only during online execution. Accordingly, the dual dynamics sample different policies over time through policy-switching. Should a stochastic policy not be needed to solve the problem, the dual dynamics converge and the state-augmented algorithm is, although viable, not needed. Moreover, the Lagrangian minimizing policies do not depend on the constraint tolerances, only the optimal dual multipliers do. Given that the dual multipliers sampled during training cover a suitable range, a single state-augmented model could easily adapt to the changes (within a reasonable range) in the constraint specifications after deployment. With conventional primal-dual training, it would be necessary to either train different models for different operation regimes or condition the model on the constraint levels during training.
One limitation of the state-augmented solution is that the choice of the uniform random sampling scheme for the dual multipliers is arbitrary and dependent on the application at hand. Second, we execute a suboptimal policy during the transient phase in the beginning before the dual multipliers reach their near-optimal values. Further research on near-optimal sampling and initialization of state-augmenting dual multipliers could significantly improve the proposed approach.
We presented network slicing as a timely and relevant application of state-augmentation. Many other wireless control and resource allocation problems of practical importance (e.g., multiple-access channels, power control, beamforming, routing and link scheduling etc.) are captured by the abstract problem formulation in (3) and can be tackled with a state-augmented primal-dual algorithm.
@@ VI Conclusion and Future Work
We formulated (Wi-Fi) network slicing as a constrained learning problem. Through numerical simulations, we illustrated that a state-augmented algorithm that minimizes an expected Lagrangian offline and updates the dual multipliers online lends itself to a practical primal-dual algorithm with feasibility and optimality guarantees for network slicing. Extension to a state-augmented RL algorithm for network slicing, exploring the impact of imperfect CSI and end-to-end evaluation on a more realistic simulation platform are interesting future research directions.